# Crosslingual Capabilities and Knowledge Barriers in Multilingual Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are typically multilingual due to pretraining on diverse multilingual corpora. But can these models relate corresponding concepts across languages, effectively being crosslingual? This study evaluates six state-of-the-art LLMs on inherently crosslingual tasks. We observe that while these models show promising surface-level crosslingual abilities on machine translation and embedding space analyses, they struggle with deeper crosslingual knowledge transfer, revealing a crosslingual knowledge barrier in both general (MMLU benchmark) and domain-specific (Harry Potter quiz) contexts. We observe that simple inference-time mitigation methods offer only limited improvement. On the other hand, we propose fine-tuning of LLMs on mixed-language data, which effectively reduces these gaps, even when using out-of-domain datasets like WikiText. Our findings suggest the need for explicit optimization to unlock the full crosslingual potential of LLMs. Our code is publicly available at https://github.com/google-research/crosslingual-knowledge-barriers.
@@ 1 Introduction
Figure 1: LLMs pretrained on internet-scale corpora containing texts in different languages are typically multilingual. While they show promising crosslingual abilities on explicit tasks like machine translation, they struggle to bridge the language gap on knowledge-intensive tasks that require implicit crosslingual correlation, revealing a crosslingual knowledge barrier.
@@
Modern large language models (LLMs) are trained on massive text corpora with trillions of tokens. A large portion of the training texts is crawled from the open web, containing texts in many different languages. As a result, many LLMs can operate in multiple languages. For example, Mistral-Large and Mixtral 8$\times$22B (Mistral, 2024) reported performance on the benchmark datasets (e.g., MMLU (Hendrycks et al., 2021), Arc Challenge (Clark et al., 2018)) in multiple languages.
@@
For humans, knowing multiple languages (multilingual) naturally implies knowing the correspondence between the words and phrases of the same meaning across those languages (crosslingual). This is because when exposed to different linguistic environments, people can develop crosslingual capabilities by grounding the languages in physical world interactions. For example, we can relate the English world “apple” to the Spanish word “manzana” because in both linguistic environments the corresponding words refer to the same fruit in the real world. On the other hand, modern LLMs are trained purely based on the statistical relations in the text corpus without any grounding in the real world. In specific tasks such as machine translation, in order to teach the models to correlate notions across different languages, it is common to train with parallel corpora — collections of pairs of texts with the same meaning but in different languages (Eisenstein, 2019). This motivates our central research question:
How well do multilingual LLMs, not explicitly  [^1] This criterion allows the case when a small amount of parallel texts accidentally crawled from the web get mixed in the pretraining dataset, as it is nearly impossible to verify. We believe such presence, if it exists at all, would have negligible impact on the model given the size of the rest of the pretraining datasets. trained on parallel corpora, exhibit crosslingual capabilities?
@@
To state the problem more precisely, we define222 We leave some of the terms mathematically vague, as long as they are not conceptually ambiguous. E.g., to measure the performance with a given correct answer, depending on the specific task format, we could either ask the model to generate the specific sequence of tokens or to rank the correct answer among multiple choices. the multilingual and crosslingual capabilities as follows. Denote an instance of a given task $\mathsf{T}$ as a tuple $(\mathcal{K},\mathcal{C},\mathcal{O})$, where $\mathcal{K}$ is the (optional) knowledge learned from training data, $\mathcal{C}$ is a context, and $\mathcal{O}$ is the correct answer. The multilingual performance on $\mathsf{T}$ measures the average performance across each language $\ell$ on an evaluation set of task instances $\{(\mathcal{K}_{\ell},\mathcal{C}_{\ell},\mathcal{O}_{\ell})\}$, where the subscript $\ell$ indicates the realization of the knowledge/context/answer in a specific language. On the other hand, the crosslingual performance on $\mathsf{T}$ measures the average performance on an evaluation set of crosslingual task instances $\{(\mathcal{K}_{\ell},\mathcal{C}_{\ell^{\prime}},\mathcal{O}_{\ell^{\prime\prime}})\}$, where $\ell,\ell^{\prime},\ell^{\prime\prime}$ can be different languages.
@@
For example, consider a task of $\mathsf{REPEAT}$. The multilingual performance simply measures the model’s capability of copying the context provided in different languages (i.e., $\mathcal{O}_{\ell}=\mathcal{C}_{\ell}$), whereas the crosslingual version of the task is equivalent to a much more challenging task of translation (i.e., $\mathcal{O}_{\ell^{\prime\prime}}=\text{Translate}_{\ell^{\prime}\Rightarrow\ell^{\prime\prime}}(\mathcal{C}_{\ell^{\prime}})$). Another example task is question-answering (QA), where the crosslingual version requires the model to apply knowledge learned from one language to compute the answer in a different language.
With those definitions, we summarize the main studies and contributions as below:
@@
Crosslingual capabilities (Section 2): We identify and formulate the question of multilingual vs crosslingual capabilities in LLMs without parallel-corpus training. Through both translation tests and embedding distance evaluations, we confirm that modern LLMs have strong crosslingual capabilities.
@@
Crosslingual knowledge barrier (Section 3): We design crosslingual QA tasks, and observe a crosslingual knowledge barrier: LLMs have a significant performance gap in QA tasks formulated in a different language from the original language in which the knowledge is learned (see Figure 1 for an illustration). Via extensive experiments across six models, including both open-source and proprietary, we confirm a systematic presence of such barriers to knowledge learned both during the pretraining and fine-tuning stages.
@@
Towards overcoming the barrier (Section 4): We propose a simple mixed-language training strategy and show that it can effectively reduce the knowledge barrier, significantly outperform other baseline methods based on prompt engineering, and further improve the few-shot learning performance. Furthermore, we show that even mixed-language training on out-of-domain data can be effective.
@@ 2 Multilingual LLMs have competitive crosslingual capabilities
In this section, we demonstrate the crosslingual capabilities of existing multilingual LLMs from two perspectives: machine translation performance (Section 2.1) and an analysis of multilingual text embeddings (Section 2.2). Specifically, our investigation focuses on five widely spoken languages: English (en), French (fr), German (de), Spanish (es), and Italian (it).
@@ 2.1 Machine translation performance
Evaluated multilingual language models. Our evaluation primarily considers six different LLMs that have exhibited strong multilingual capabilities. These include four open-source models: Llama2-7B, Llama2-13B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Llama3-8B (Meta, 2024), and two proprietary models: GPT-3.5 and GPT-4 (Achiam et al., 2023). Section B.1 provides HuggingFace links or endpoint specifications for all evaluated models. To perform machine translation tasks with the open-source models, we use the prompting format proposed by Xu et al. (2024). For proprietary models, we use the prompting template suggested on their official webpages. [^3] https://platform.openai.com/examples/default-translation
Strong baselines. For reference we report two strong baselines: 1) the NLLB-3.3B model, the largest supervised encoder-decoder translation model from the NLLB family (Costa-jussà et al., 2022), which was trained on parallel corpus from various sources for 204 languages; and 2) the Google Translate API.
Multilingual LLMs achieve competitive performance in machine translation. We evaluate translation performance on a subset of $100$ examples from the FLoRes-101 benchmark (Goyal et al., 2022) for two directions per language: translating from English to the target language (en $\rightarrow$ X), and from the target language to English (X $\rightarrow$ en). We report the COMET score computed using the COMET-22 model (Rei et al., 2020), a metric designed to predict human judgments of machine translation quality, following previous work (Zhu et al., 2024; He et al., 2024; Xu et al., 2024).
As shown in Table 1, even though the evaluated multilingual LLMs are not directly trained on parallel corpora, their translation ability is quite competitive when compared to translation models explicitly trained on parallel corpora or industrial-grade translation APIs. Notably, these models generally perform better when translating non-English text into English, but worse in the opposite direction, potentially suggesting that they are more proficient with the English translations. Those results are consistent with previous papers that focus on improving machine translation with pretrained LLMs (Zhu et al., 2024; He et al., 2024; Xu et al., 2024). However, our evaluation has a different goal of comprehensively examining LLMs’ crosslingual capabilities beyond translation.
Table 1: COMET scores for machine translation tasks evaluated on a FloRes-101 subset benchmark using multilingual LLMs (Llama2-7B, Mistral-7B, Llama2-13B, Llama3-8B, GPT-3.5, and GPT-4), models trained on parallel corpora (NLLB-3.3B), and an industrial-grade translation API (Google Translate). Multilingual LLMs achieve competitive translation performance against dedicated translation models and the translation API.
Llama2-7B $83.93$ $85.82$ $85.93$ $84.67$ $85.09$ $86.94$ $88.70$ $85.52$ $86.73$ $86.97$
Mistral-7B $77.94$ $78.23$ $79.13$ $78.13$ $78.36$ $87.48$ $88.38$ $85.64$ $86.75$ $87.06$
Llama3-8B $83.24$ $86.56$ $84.99$ $85.98$ $85.19$ $89.22$ $88.84$ $87.04$ $87.78$ $88.22$
GPT-3.5 $87.53$ $86.97$ $86.40$ $86.46$ $86.84$ $89.14$ $89.42$ $87.47$ $88.41$ $88.61$
GPT-4 $87.29$ $87.56$ $86.49$ $86.64$ $87.00$ $89.27$ $89.66$ $87.53$ $88.42$ $88.72$
NLLB-3.3B $87.33$ $87.44$ $86.88$ $88.26$ $87.48$ $79.65$ $87.44$ $85.64$ $84.13$ $84.72$
Google Translate $89.13$ $88.77$ $87.25$ $89.21$ $88.59$ $89.96$ $90.15$ $87.47$ $88.80$ $89.10$
@@ 2.2 Embedding of mixed translated sentences
We further investigate the explicit crosslingual ability of multilingual LLMs by probing their text embeddings. Specifically, we aim to verify whether the embeddings for a given text in English are similar to the embeddings when some words are presented in different languages. We focus primarily on open-source models due to the cost associated with querying embeddings from proprietary models.
Experimental setup. We randomly sample $1,000$ examples from the WikiText-103 corpus (Merity et al., 2017). For each example, we create two versions: (1) The original text in English; (2) Mixed-language-translated: for each word, with a probability of $p=0.8$ it is unchanged; and with a probability of $1-p$, the word is (independently) translated, using Google Translate API, into a random language selected from the set {en, fr, de, es, it}. In other words, each word has a $0.16$ probability of being translated into a non-English language. We then obtain sentence embeddings from the LLM for both versions of each example.
@@
To establish baselines for comparison, we consider two scenarios representing an “upper bound” on the distance when perturbations are unrelated to the original content: (1) Random Token Replacement: with a probability of $p=0.16$, each token is replaced with a random different token from the vocabulary; and (2) Random Token Dropout: with a probability of $p=0.16$, a token is completely masked out by disallowing any attention to it.
@@
The original embedding size is 4096. To visualize and compare the embeddings, we perform non-linear dimensionality reduction (McInnes et al., 2018), reducing each embedding to a 2D vector. For each example, we calculate the per-coordinate distance between the 2D embedding vectors of the original English text and the mixed-language-translated text. We then visualize these distance values and compare them to the baseline scenarios calculated in the same way, as shown in Figure 2.
@@
Embeddings of English and mixed-language-translated text are similar. As shown, the embeddings of the original text and the mixed-translated text are quite close, with the difference vectors scattered around the origin. To further quantify this observation, we perform a two-sample statistical test on the following two distributions: (1) the cosine similarities between embeddings of original sentences and the embeddings of corresponding mixed-translated sentences, and (2) the cosine similarities between embeddings of original sentences and embeddings of corresponding sentences with random token replacement. The $p$-value of the test is smaller than $0.05$ for all models, suggesting that these two distributions are significantly different. In other words, replacing English words with translated words is significantly different from replacing English words with random tokens, which implies the explicit crosslingual capabilities of the multilingual LLMs.
3 Crosslingual knowledge barrier in general and domain-specific knowledge
While multilingual LLMs have demonstrated impressive explicit crosslingual abilities, as we will show in this section, they also struggle to seamlessly bridge the language gap when faced with tasks demanding implicit crosslingual knowledge transfer. We term this phenomenon the crosslingual knowledge barrier. In the following, we demonstrate the presence of such barriers for both general knowledge (Section 3.1) and domain-specific knowledge (Section 3.2).
3.1 Crosslingual knowledge barrier in general knowledge
Monolingual evaluation is inadequate for assessing crosslingual abilities. Previous studies have evaluated open-domain Multiple Choice Question (MCQ) tasks on general knowledge in multilingual settings. For instance, Mistral series models (Mistral, 2024) were benchmarked on translated versions of the Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) dataset in languages such as French (fr), German (de), Spanish (es), and Italian (it), separately. We call such monolingual evaluation setups as “full-translation” (i.e., the full English MCQ dataset is translated into another language).
While such results indicate multilingual proficiency, they are insufficient to show crosslingual proficiency. The core issues is that the relevant general knowledge might be present in each of the evaluated languages in the pretraining dataset, so the LLMs could answer the translated questions based on knowledge learned in each individual language without invoking any crosslingual capabilities. Such possibility is difficult to verify, as the pretraining datasets for most existing LLMs are undisclosed.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Monolingual evaluation of six LLMs on MMLU (fully translated for non-English languages), which is additionally split into four domains (STEM, Social Science, Humanities, Others) based on its subject categories. LLMs consistently perform better at answering multi-choice questions in English than in other languages.
Mixed-language evaluation. To directly evaluate crosslingual capabilities of LLMs on general knowledge MCQ tasks, we suggest adopting an inherent crosslingual interaction approach through mixed-language MCQ formats. Specifically, we propose the following formats purposefully designed to be novel compositions unlikely to have been encountered during pretraining (examples can be found in Figure 4):
  • Mixup translation: translating the question and all options into 5 different languages, with the language assignments randomly determined from the set {en, fr, de, es, it}.
  • Question translation: translating the question into a non-English language.
  • Options translation: translating all options into a non-English language.
  • Question+GT-option translation: translating both the question and the ground truth option into a non-English language, while keeping the remaining options in English.
  • GT-option translation: translating the ground truth option into a non-English language, while keeping the question and the rest of the options in English.
  • One-wrong-option translation: randomly selecting one incorrect option and translating it into a non-English language.

### v2

@@ Abstract
Large language models (LLMs) are typically multilingual due to pretraining on diverse multilingual corpora. But can these models relate corresponding concepts across languages, i.e., be crosslingual? This study evaluates state-of-the-art LLMs on inherently crosslingual tasks. We observe that while these models show promising surface-level crosslingual abilities on machine translation and embedding space analyses, they struggle with deeper crosslingual knowledge transfer, revealing a crosslingual knowledge barrier in both general (MMLU benchmark) and domain-specific (Harry Potter quiz and TOFU benchmark) contexts. Since simple inference-time mitigation methods offer only limited improvement, we propose fine-tuning of LLMs on mixed-language data, which effectively reduces these gaps, even when using out-of-domain datasets like WikiText. Our findings suggest the need for explicit optimization to unlock the full crosslingual potential of LLMs.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: While multilingual LLMs show promising crosslingual abilities on explicit tasks like machine translation where the source text is provided in the context, they struggle to bridge the language gap on knowledge-intensive tasks that require implicit crosslingual correlation of parametric knowledge, revealing a crosslingual knowledge barrier. Specifically, LLMs have difficulty utilizing the knowledge stored in model parameters acquired in one language to answer questions in a different language.
@@
Modern LLMs are trained on massive text corpora with trillions of tokens. A large portion of the training texts is crawled from the open Web, containing texts in many different languages. As a result, many LLMs can operate in multiple languages. For example, Mistral models (Mistral, 2024) reported performance on the benchmark datasets in multiple languages (e.g., MMLU (Hendrycks et al., 2021), Arc Challenge (Clark et al., 2018)).
@@
For humans, knowing multiple languages (multilinguality) naturally implies knowing the correspondence between the words / phrases of the same meaning across those languages (crosslinguality). When exposed to different linguistic environments, people can develop crosslingual capabilities by grounding in physical world interactions. For example, we can relate the English world “apple” to the Spanish word “manzana” because in both linguistic environments the corresponding words refer to the same fruit in the real world. On the other hand, modern LLMs are trained purely based on the statistical relations in the text corpus without any grounding in the real world. In specific tasks such as machine translation, in order to teach the models to correlate notions across different languages, it is common to train with parallel corpora — collections of pairs of texts with the same meaning but in different languages (Eisenstein, 2019). However, as the training process of most LLMs is often unknown, it is difficult to ascertain whether or to what extent crosslingual supervision like parallel corpora were employed. This is particularly relevant for models that naturally perform well in multiple languages due to their massive web training data, even though they might not be explicitly advertised to target multilingual capabilities. [^1] Recent efforts mine massive parallel texts from the web (Schwenk et al., 2021), which may have been used in the pre-training datasets of some LLMs, particularly those designed with multilingual capabilities. This ambiguity motivates our central research question: How well do multilingual LLMs exhibit crosslingual capabilities?
@@
To state the problem more precisely, we define the multilingual and crosslingual capabilities as follows. Denote an instance of a given task $\mathsf{T}$ as a tuple $(\mathcal{K},\mathcal{C},\mathcal{O})$, where $\mathcal{K}$ is the (optional) knowledge learned from training data, $\mathcal{C}$ is a context, and $\mathcal{O}$ is the correct answer. The multilingual performance on $\mathsf{T}$ measures the average performance across each language $\ell$ on an evaluation set $\{(\mathcal{K}_{\ell},\mathcal{C}_{\ell},\mathcal{O}_{\ell})\}$ of task instances. On the other hand, the crosslingual performance on $\mathsf{T}$ measures the average performance on an evaluation set $\{(\mathcal{K}_{\ell},\mathcal{C}_{\ell^{\prime}},\mathcal{O}_{\ell^{\prime\prime}})\}$ of cross-linguall task instances, where $\ell,\ell^{\prime},\ell^{\prime\prime}$ can be different languages.
@@
For example, consider a task of REPEAT. The multilingual performance simply measures the model’s capability of copying the context provided in different languages (i.e., $\mathcal{O}_{\ell}=\mathcal{C}_{\ell}$), whereas the crosslingual version of the task is equivalent to translation task, i.e., $\mathcal{O}_{\ell^{\prime\prime}}=\text{Translate}_{\ell^{\prime}\Rightarrow\ell^{\prime\prime}}(\mathcal{C}_{\ell^{\prime}})$. Notably, in translation, the source text invoking crosslingual ability is explicitly provided in context. A more challenging scenario arises when the task involves implicit crosslingual knowledge reasoning. In particular, we consider crosslingual knowledge question-answering (QA) task, which requires LLMs to implicitly retrieve and utilize parametic knowledge $\mathcal{K}_{\ell}$ learned from one language $\ell$ to answer questions in a different language $\ell^{\prime}$.
With those definitions, we summarize our contributions:
@@
Crosslingual capabilities (§ 2): We formulate the question of multilingual vs crosslingual capabilities in LLMs. Through both translation tests (§ 2.1) and embedding distance evaluations (§ 2.2), we confirm that modern LLMs have competitive crosslingual capabilities.
@@
Crosslingual knowledge barrier (§ 3): We design crosslingual QA tasks, and observe a crosslingual knowledge barrier on 15 models, 16 languages and 3 datasets: LLMs have a significant performance gap on QA tasks formulated in a different language from the original language in which the knowledge is learned (see Fig. 1). Via extensive experiments, we confirm the presence of such barriers to knowledge learned both during the pretraining (§ 3.1) and fine-tuning (§ 3.2) stages. As far as we know, our study is the first to systematically identify a cross-lingual barrier for using parametric knowledge in modern LLMs, covering both general and domain-specific knowledge.
@@
Towards overcoming the barrier (§ 4): We propose a mixed-language training strategy (§ 4.2) and show that it can effectively reduce the knowledge barrier, outperform other baseline methods based on prompt engineering (§ 4.1), and further improve the few-shot learning performance. Furthermore, we show that 1) even mixed-language fine-tuning on out-of-domain corpus can be effective; 2) it can enhance model performance on out-of-distribution languages that were not included during mixed-language fine-tuning (§ 4.2).
@@ 2 Multilingual LLMs have competitive crosslingual capabilities
We demonstrate crosslingual capabilities of multilingual LLMs through machine translation performance (§ 2.1) and multilingual text embeddings analysis(§ 2.2).
Evaluation focus. (1) Languages: We focus on five widely spoken languages: English (en), French (fr), German (de), Spanish (es), and Italian (it). Since our crosslingual study relies on the model being multilingual (i.e., that it already knows the languages well), we chose to evaluate these languages, as explicitly mentioned in the reports of some LLMs (Mistral, 2024). (2) LLMs: We mainly focus on six popular LLMs, including four open-source models: Llama2-7B, Llama2-13B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Llama3-8B (Meta, 2024), and two proprietary models: GPT-3.5 and GPT-4 (Achiam et al., 2023). § D.1 provides specifications for those LLMs.
@@ 2.1 Machine translation performance
Setup. To perform machine translation tasks with the open-source LLMs, we use the prompting format proposed by Xu et al. (2024). For proprietary LLMs, we use the prompting template suggested on their official webpages. [^2] https://platform.openai.com/examples/default-translation For reference we report two strong baselines: 1) NLLB-3.3B, the largest supervised encoder-decoder translation model from the NLLB family (Costa-jussà et al., 2022) trained on parallel corpus for 204 languages; and 2) Google Translate API. We report translation performance measured by the COMET score (Rei et al., 2020), a metric to predict human judgments of machine translation quality, on FLoRes-101 benchmark (Goyal et al., 2022) for two directions per language: en $\rightarrow$ X and X $\rightarrow$ en.
Table 1: COMET scores for machine translation tasks evaluated on FloRes-101 benchmark. The score for X is averaged over {fr,de, es, it}.
en $\rightarrow$ X X $\rightarrow$ en
Llama2-7B $84.04$ $86.91$
Llama2-13B $77.81$ $87.50$
Llama3-8B $79.32$ $87.65$
Mistral-7B $77.83$ $86.90$
Qwen2.5-7B $86.37$ $88.27$
Llama-3.1-8B $83.94$ $87.39$
GPT-3.5 $86.84$ $88.61$
GPT4 $87.07$ $88.71$
NLLB-3.3B $87.48$ $84.72$
Google Translate $88.80$ $89.01$
Figure 2: Embeddings of En text and mixed-language-translated text are more closely aligned than baselines. The ellipses represent the covariance confidence intervals.
Multilingual LLMs achieve competitive performance in machine translation. As shown in Tb. 1 (more results deferred to Tb. 7), the translation ability of evaluated multilingual LLMs is quite competitive when compared to translation models explicitly trained on parallel corpora or industrial-grade translation APIs (e.g., the gap is within 2.11 COMET score for X $\rightarrow$ en translation.) Notably, these models generally perform better when translating X $\rightarrow$ en, but worse in the opposite direction, potentially suggesting that they are more proficient with English translations. These results are consistent with previous papers that focus on improving machine translation with pretrained LLMs (Zhu et al., 2024; He et al., 2024; Xu et al., 2024). However, as we will show in § 3, our study focuses on crosslingual transferability of knowledge implicitly store in model weights, beyond the direct translation task where the source text is explicitly provided in context.
@@ 2.2 Embedding of mixed translated sentences
We further investigate the explicit crosslingual ability of multilingual LLMs by probing their text embeddings. Specifically, we verify whether the embeddings for a given text in English are similar to the embeddings when some words are presented in different languages. The embedding is a vector as the average of the last layer’s activations across all tokens in the sentence (BehnamGhader et al., 2024).
@@
Setup. We randomly sample $1k$ examples from the WikiText-103 corpus (Merity et al., 2017), creating two versions for each: (1) The original text in English; (2) Mixed-language-translated: for each word, with a probability of $0.2$ it is (independently) translated into a random language selected from {en, fr, de, es, it} using Google Translate API. That is, each word has a $p=0.16$ probability of being translated into a non-English language. To establish baselines for comparison, we consider two scenarios representing an “upper bound” on the distance when perturbations are unrelated to the original content: (1) Random Token Replacement: with a probability of $p=0.16$, each token is replaced with a random different token from the vocabulary space of the tokenizer; and (2) Random Token Dropout: with a probability of $p=0.16$, a token is completely masked out by disallowing any attention to it.
@@
To visualize and compare embeddings, we reduce the original 4096-dimensional vectors to 2D using non-linear dimensionality reduction. We then calculate and visualize the per-coordinate distances between 2D embeddings of original English text and mixed-language translations, comparing these to baseline scenarios.
@@
As shown in Fig. 2 (more results in Fig. 11), embeddings of English and mixed-language-translated text exhibit greater similarity compared to the baselines, with difference vectors clustered near the origin. To quantify this, we conducted a two-sample statistical test comparing cosine similarities between: (1) original and mixed-translated sentence embeddings, and (2) original and random-token-replaced sentence embeddings. The resulting $p$-value ($<0.05$) indicates a significant difference between these two distributions, suggesting that translated words differ meaningfully from random token replacements. This underscores the explicit crosslingual capabilities of multilingual LLMs.
3 Identify crosslingual knowledge barrier